how do I find the hypotenuse of a triangle if side c=12 and side b=8?
Answer by stanbon(75887)

how do I find the hypotenuse of a triangle if side c=12 and side b=8?
--------------
Use Pythagoras to solve for the hypotenuse:
h^2 = 12^2 + 8^2
h^2 = 144 + 64
h^2 = 208
h= 14.42
==============
Cheers,
Stan H.